OpenStudy (anonymous):

what is the molality of a water solution that has a boiling point of 102.00 degrees celcious? water has a boiling point of 0.52 degrees celcious/m.

OpenStudy (anonymous):

\[\Delta Tb = 102 - 100 = 2\] \[\Delta Tb = m \times Kb\] 2 = m x 0.52 m = 2/0.52 = 3.84 molal

OpenStudy (anonymous):

thaaank youuu

OpenStudy (anonymous):

ur welcome